OpenStudy (anonymous):

factorise the eq : x^2 - 3x +9/4

OpenStudy (amistre64):

notice that its a perfect square?

OpenStudy (anonymous):

what two number sum to -3 and multiplies to 9/4

OpenStudy (anonymous):

\[(4x ^{2}-6x+9)/4= (2x-3)(2x-3)/4\]

OpenStudy (anonymous):

nonononon. the constant c = 9/4. Not the whole thing over 4.

OpenStudy (anonymous):

yes it is i multiplied the numerator by 4 as a sort of LCM hope u understood

OpenStudy (anonymous):

can't it be solved without using an LCM?

OpenStudy (anonymous):

ya

OpenStudy (amistre64):

try completeing the square on it :)

OpenStudy (anonymous):

its a perfect square, you can'r, and don't need to complete the square!

OpenStudy (anonymous):

yes it can be done like dis.....\[x ^{2}-2*(3/2)*x+(3/2)^{2}=(x-3/2)^{?}\]

OpenStudy (anonymous):

I asked you again , what two number sum to -3 and multiplies to 9/4

OpenStudy (amistre64):

since its a perfect square; then just compact it then

OpenStudy (anonymous):

-3/2 and -3/2 add up to -3 and product is 9/4

OpenStudy (amistre64):

(x + b/2)^2

OpenStudy (anonymous):

\[x^2 -3x + \frac{9}{4}\]

OpenStudy (anonymous):

so it would be (x-3/2)^2 ? Yes, thank you agdgdgdgwngo.

OpenStudy (anonymous):

so \[(x-\frac{3}{2})^2\]

OpenStudy (anonymous):

4x^2-12x+9=4x^2-6x-6x+9=2x(2x-3)-3(2x-3)=(2x-3)^2

OpenStudy (anonymous):

Thank you, imranmeah91.